# On torsion of non-acyclic cellular chain complexes of even manifolds in a unique factorisation monoid

### v1

@@ Abstract.
Let $\mathcal{M}_{2n}^{\mathrm{Diff},\mathrm{hc}}$ be a multiplicative factorisation monoid over highly connected differentiable closed connected oriented manifolds. Any $2n$-dimensional manifold $W_{p}^{2n}$ from $\mathcal{M}_{2n}^{\mathrm{Diff},\mathrm{hc}}$ admits a unique connected sum decomposition into manifolds that cannot be decomposed any further. By using this decomposition, we prove that Reidemeister-Franz torsion of $W_{p}^{2n}$ can be written as the product of Reidemeister-Franz torsions of the manifolds in the decomposition without the corrective term.
@@ 1. Introduction
All manifolds are considered as non-empty, closed, connected and oriented. An $n$-manifold $M^{n}$ is called highly connected if $\pi_{i}(M^{n})=0$ for $i=0,\ldots,\lfloor n/2\rfloor-1.$ Let us denote the diffeomorphism classes of $n$-dimensional differentiable manifolds by $\mathcal{M}_{n}^{\mathrm{Diff}}.$ Thus, the diffeomorphism classes of $n$-dimensional highly connected differentiable manifolds is given as follows
$\mathcal{M}_{n}^{\mathrm{Diff},\mathrm{hc}}=\{M^{n}\in\mathcal{M}_{n}^{\mathrm{Diff}}|M^{n}\;\mathrm{is\;highly\;connected}\}.$
For $n\in\mathbb{N}$, let $M^{n},N^{n}\in\mathcal{M}_{n}^{\mathrm{Diff}}.$ Given an orientation-preserving smooth embedding $\varphi:\overline{\mathbb{D}^{n}}\rightarrow M^{n}$ and given an orientation-reversing smooth embedding $\varrho:\overline{\mathbb{D}^{n}}\rightarrow N^{n},$ the connected sum of $M^{n}$ and $N^{n}$ is defined as
$M^{n}\#N^{n}=((M^{n}\setminus\varphi({\mathbb{D}^{n}})\sqcup(N^{n}\setminus\varrho({\mathbb{D}^{n}}))/\varphi(C)=\varrho(C)\mathrm{\ for\ all\ }C\in\mathbb{S}^{n-1}.$
The diffeomorphism type of the connected sum of two differentiable manifolds is independent of the choice of embedding, [19, Theorem 2.7.4]. Hence, by [1], $\mathcal{M}_{n}^{\mathrm{Diff}}$ and its subset $\mathcal{M}_{n}^{\mathrm{Diff},\mathrm{hc}}$ are abelian monoids (written multiplicatively) under connected sum operation.
@@
The Reidemeister-Franz torsion is a well-known topological invariant [2, 6, 8, 13]. It has been instrumental in disproving Hauptvermutung. Moreover, Milnor described Reidemeister-Franz torsion with the Alexander polynomial [9, 10].
Assume that $W^{2n}\in\mathcal{M}_{2n}^{\mathrm{Diff},\mathrm{hc}}$ is decomposed into two closed, oriented manifolds $M_{L}^{2n}$ and $M_{R}^{2n}.$ In [10, Theorem 3.2], Milnor showed that Reidemeister-Franz torsion acts multiplicatively with respect to such gluings. Namely, the torsion of $W^{2n}$ is the product of the torsions of $M_{L}^{2n},$ $M_{R}^{2n},$ and the torsion of (2n1)(2n-1)-sphere $\mathbb{S}^{2n-1}$ times a corrective term $\mathbb{T}_{RF}\left(\mathcal{H}_{\ast}\right)$ coming from homologies of the chain complexes of the cell-decompositions of manifolds. By [10, Theorem 3.1], if the chain complex of the cell-decomposition of the manifold is acylic, then the corrective term becomes $1.$ Without the acyclicity assumption, while there are lots of examples of computing the Reidemeister-Franz torsions of $2$ and $3$-dimensional manifolds [4, 5, 12, 15, 16], there are not so many examples of computations for higher dimensional manifolds. Our main goal is to give an example of a class of higher-dimensional manifolds that is closed under the connected sums for which corrective terms become 1 without the assumption of acyclicity.
@@ 2.3. The Reidemeister-Franz torsion of a manifold
From the lemma above, we can conclude that the Reidemeister-Franz torsion of $M^{n}$ is well-defined. So we denote by $\mathbb{T}_{RF}(M^{n},\{\mathbf{h}_{p}\}_{0}^{n})$ the Reidemeister-Franz torsion of $M^{n}$ in the basis $\mathbf{h}_{p}$ of $H_{p}(M^{n})$ for $p\in\{1,\ldots,n\}$.
@@ 3. Main results
As a warm up, we are going to start this section by calculating the torsion of the closed unit ball $\overline{\mathbb{D}^{2n}}.$ Next, to prove Theorem 1.0.5, we need to calculate the Reidemeister-Franz torsion of ${M^{2n}}-{\mathbb{D}^{2n}}$ in terms of the torsion of $M^{2n}$} (Theorem 3.3.1). Later, we give a formula to calculate the torsion of $W^{2n}=M_{L}^{2n}\#M_{R}^{2n}$ in terms of torsions of $M_{L}^{2n}$ and $M_{R}^{2n}$ (Theorem 3.2.2). Therefore, these calculations form a template for the homological algebraic calculations that we need for the proof of Theorem 1.0.5.
@@ 3.2. The Reidemeister-Franz torsion of $W^{2n}=M_{L}^{2n}\#M_{R}^{2n}$
If we consider equation (3.2.5) for $C_{1}(\mathcal{H}_{\ast})=H_{0}({M_{L}^{2n}}-{\mathbb{D}}^{2n})\oplus H_{0}({M_{R}^{2n}}-{\mathbb{D}}^{2n}),$ then we have
@@
$H_{2n-1}(\mathbb{S}^{2n-1})\overset{\imath_{2n-1}^{\ast}}{\rightarrow}H_{2n-1}({M_{L}^{2n}}-{\mathbb{D}}^{2n})\oplus H_{2n-1}({M_{R}^{2n}}-{\mathbb{D}}^{2n})\overset{\pi_{2n-1}^{\ast}}{\rightarrow}H_{2n-1}(W^{2n}).$
@@
(3.2.19) $[\mathbf{h}_{6n+1}\rightarrow\mathbf{h}^{\prime}_{6n+1}]=1,$
$[\mathbf{h}_{6n+2}\rightarrow\mathbf{h}^{\prime}_{6n+2}]=1.$
@@
If we consider equations (3.2.7), (3.2.9), (3.2.11), (3.2), (3.2.16), (3.2.18), (3.2.19), (3.2), (3.2.22), and (3.2.24), then we show that the corrective term satisfies the following equation
@@ 3.3. The Reidemeister-Franz torsion of ${M^{2n}}-{\mathbb{D}^{2n}}$
$\mathbf{h}^{\prime}_{1}=\left\{\mathbf{h}^{\mathrm{Im}\,\imath_{0}^{\ast}},s_{1}(\mathrm{Im}\,\pi_{0}^{\ast})\right\}$
@@
in the previous step. It follows from equation (3.3.6) that $s_{2}(\mathrm{Im}(\imath_{0}^{\ast}))$ is the obtained basis $\mathbf{h}^{\prime}_{2}$ of $C_{2}(\mathcal{H}_{\ast}).$ If we take the initial basis $\mathbf{h}_{2}$ (namely, $\mathbf{h}_{0}^{\mathbb{S}^{2n-1}}$) of $C_{2}(\mathcal{H}_{\ast})$ as $s_{2}(\mathrm{Im}(\imath_{0}^{\ast})),$ then we obtain
@@
By the exactness of the long-exact sequence $\mathcal{H}_{\ast}$, Lemma 3.2.1, and the First Isomorphism Theorem, we obtain the followings for each $i\in\{1,\ldots,2n-1\}$
@@
For each $i\in\{1,\ldots,2n-1\},$ by using the isomorphism $H_{i}({M^{2n}}-{\mathbb{D}^{2n}})\overset{\pi^{\ast}_{i}}{\cong}H_{i}(M^{2n})$ and given basis $\mathbf{h}^{{M^{2n}}-{\mathbb{D}^{2n}}}_{i}$ of $H_{i}({M^{2n}}-{\mathbb{D}^{2n}}),$ we can consider the basis of $\mathbf{h}^{M^{2n}}_{i}$ of $H_{i}(M^{2n})$ as $\pi^{\ast}_{i}(\mathbf{h}^{{M^{2n}}-{\mathbb{D}^{2n}}}_{i}).$ As $\imath_{i}^{\ast}$ is zero map, $\mathrm{Im}\,\imath_{i}^{\ast}=\{0\}$. Since $\pi^{\ast}_{i}$} is an isomorphism, its inverse can be considered as the section $s_{{}_{i}}$. As in the proof of Theorem 3.2.2, we obtain
@@
(3.4.2) $\mathbb{T}_{RF}(W_{i}^{2n},\{\mathbf{h}_{\nu}^{W_{i}^{2n}}\}_{0}^{2n})$ $=$ $\mathbb{T}_{RF}({W_{i-1}^{2n}}-{\mathbb{D}_{i}^{2n}},\{\mathbf{h}_{\nu}^{{W_{i-1}^{2n}}-{\mathbb{D}_{i}^{2n}}}\}_{0}^{2n})$
@@
and their associated Mayer-Vietoris long exact sequences as in Proposition 3.3.1. In Proposition 3.3.1, $\mathbf{h}_{\nu}^{W_{i-1}^{2n}}$ and $\mathbf{h}_{\nu}^{{M_{i+1}^{2n}}-{\mathbb{D}_{i}^{2n}}}$ are any given homology bases. So we can take these bases as above which is satisfying the equation (3.4.2). Since it is arbitrarily given basis in Proposition 3.3.1, we can respectively choose the same bases $\mathbf{h}_{0}^{\mathbb{S}^{2n-1}}$ and $\mathbf{h}_{2n-1}^{\mathbb{S}^{2n-1}}=\delta_{2n}(\mathbf{h}_{2n}^{W^{2n}})$ of $H_{0}(\mathbb{S}^{2n-1})$ and $H_{2n-1}(\mathbb{S}^{2n-1})$ for both sequences (3.4.3) and (3.4.4). Hence, for the basis $\mathbf{h}_{0}^{\overline{\mathbb{D}_{i}^{2n}}}=f^{i}_{\ast}(\varphi_{0}(\mathbf{c}_{0}))$ of $H_{0}(\overline{\mathbb{D}_{i}^{2n}}),$ there exist respectively bases $\mathbf{h}_{\nu}^{W_{i-1}^{2n}}$ and $\mathbf{h}_{\nu}^{M_{i+1}^{2n}}$ of $H_{\nu}^{W_{i-1}^{2n}}$ and $H_{\nu}^{M_{i+1}^{2n}}$ such that the following formulas hold
@@
By combining equations (3.4.2), (3.4.5), and (3.4.6), we obtain the Reidemeister-Franz torsion of $W_{i}^{2n}$ with untwisted $\mathbb{R}$-coefficients in these homology bases as follows

### v2

# Multiplicativity of Reidemister-Franz Torsion for Even Manifolds

@@ Abstract.
We study Reidemeister-Franz torsion for non-acyclic cellular chain complexes arising from closed, oriented, highly connected even dimensional manifolds. The monoid of such manifolds under connected sum admits a unique factorisation into indecomposable elements. Using this factorisation, we prove that the Reidemeister-Franz torsion of an even-dimensional manifold decomposes multiplicatively as the product of the torsions of its prime factors without any corrective term.
@@ 1. Introduction
All manifolds considered in this paper are assumed to be non-empty, closed, connected and oriented. An $n$-manifold $M^{n}$ is called highly connected if
$\pi_{i}(M^{n})=0$
for $i=0,\ldots,\lfloor n/2\rfloor-1.$ We denote by $\mathcal{M}_{n}^{\mathrm{Diff}}$ the set of the diffeomorphism classes of $n$-dimensional differentiable manifolds, and by $\mathcal{M}_{n}^{\mathrm{Diff},\mathrm{hc}}$ the subset consisting of highly connected manifolds.
For $n\in\mathbb{N}$, let $M^{n},N^{n}\in\mathcal{M}_{n}^{\mathrm{Diff}}.$ Given an orientation-preserving smooth embedding $\varphi:\overline{\mathbb{D}^{n}}\rightarrow M^{n}$ and an orientation-reversing smooth embedding $\varrho:\overline{\mathbb{D}^{n}}\rightarrow N^{n},$ the connected sum of $M^{n}$ and $N^{n}$ is defined by
$M^{n}\#N^{n}=((M^{n}\setminus\varphi({\mathbb{D}^{n}})\sqcup(N^{n}\setminus\varrho({\mathbb{D}^{n}}))/\sim,$
where the equivalence relation $\sim$ identifies $\varphi(x)\sim\varrho(x)\mathrm{\ for\ all\ }x\in\mathbb{S}^{n-1}=\partial\overline{\mathbb{D}^{n}}.$ The diffeomorphism type of the connected sum of two differentiable manifolds is independent of the choice of embedding, [19, Theorem 2.7.4]. Hence, by [1], $\mathcal{M}_{n}^{\mathrm{Diff}}$ and its subset $\mathcal{M}_{n}^{\mathrm{Diff},\mathrm{hc}}$ are abelian monoids (written multiplicatively) under connected sum operation.
@@
Reidemeister–Franz torsion is a classical topological invariant that encodes geometric and combinatorial information not detected by homology alone [2, 6, 8, 13]. Since its introduction, it has played an important role in manifold theory and related areas, most notably in the disproof of the Hauptvermutung [9, 10]. Subsequent developments, particularly those due to Milnor, established a precise relationship between Reidemeister–Franz torsion and algebraic invariants such as the Alexander polynomial, and clarified its behavior under fundamental topological constructions.
In recent years, there has been growing interest in the applied and computational aspects of topological invariants. Because of this, understanding how Reidemeister–Franz torsion behaves under decomposition and reconstruction of manifolds is of central importance. When a closed, oriented $2n$-dimensional manifold $W^{2n}\in\mathcal{M}_{2n}^{\mathrm{Diff},\mathrm{hc}}$ is obtained by gluing simpler manifolds $M_{L}^{2n}$ and $M_{R}^{2n}$, an explicit multiplicative formula for the torsion allows one to reduce global computations to local data. Moreover, the torsion of $W^{2n}$ is the product of the torsions of $M_{L}^{2n},$ $M_{R}^{2n},$ and the torsion of (2n1)(2n-1)-sphere $\mathbb{S}^{2n-1}$ times a corrective term $\mathbb{T}_{RF}\left(\mathcal{H}_{\ast}\right)$ coming from homologies of the chain complexes of the cell-decompositions of manifolds. Due to Milnor [10, Theorem 3.1], under the acyclicity assumption on the chain complexes of the cell-decompositions of such manifolds, the Reidemeister–Franz torsion is multiplicative up to the corrective term, which becomes trivial in the acyclic case.
There exist numerous explicit computations in low-dimensional settings [4, 5, 12, 15, 16]. However, little is known about Reidemeister–Franz torsion for higher-dimensional manifolds, especially when the acyclicity condition is not satisfied. The lack of such computations limits the applicability of torsion, where acyclicity cannot always be assumed and corrective terms must be carefully controlled.
The aim of this paper is to address this gap by presenting a class of higher-dimensional closed manifolds for which the multiplicativity of Reidemeister–Franz torsion holds without the assumption of acyclicity. We show that, for such manifolds, the corrective term in the multiplicative formula is trivial.
@@ 2.3. The Reidemeister-Franz torsion of a manifold
From Lemma 2.3.2, we can conclude that the Reidemeister-Franz torsion of $M^{n}$ is well-defined. So we denote by $\mathbb{T}_{RF}(M^{n},\{\mathbf{h}_{p}\}_{0}^{n})$ the Reidemeister-Franz torsion of $M^{n}$ in the basis $\mathbf{h}_{p}$ of $H_{p}(M^{n})$ for $p\in\{1,\ldots,n\}$.
@@ 3. Main results
As a warm up, we are going to start this section by calculating the torsion of the closed unit ball $\overline{\mathbb{D}^{2n}}.$ Next, to prove Theorem 1.0.5, we need to calculate the Reidemeister-Franz torsion of $W^{2n}=M_{L}^{2n}\#M_{R}^{2n}$ in terms of torsions of $M_{L}^{2n}-{\mathbb{D}^{2n}}$ and $M_{R}^{2n}-{\mathbb{D}^{2n}}$ (Theorem 3.2.2). Later, we give a formula to calculate the torsion of ${M^{2n}}-{\mathbb{D}^{2n}}$ in terms of the torsion of $M^{2n}$} (Theorem 3.3.1). Therefore, these calculations form a template for the homological algebraic calculations that we need for the proof of Theorem 1.0.5.
@@ 3.2. The Reidemeister-Franz torsion of $W^{2n}=M_{L}^{2n}\#M_{R}^{2n}$
If we consider equation (3.2.5) for the space $C_{1}(\mathcal{H}_{\ast})=H_{0}({M_{L}^{2n}}-{\mathbb{D}}^{2n})\oplus H_{0}({M_{R}^{2n}}-{\mathbb{D}}^{2n}),$ then we have
@@
$H_{2n-1}(\mathbb{S}^{2n-1})\overset{\imath_{2n-1}^{\ast}}{\rightarrow}H_{2n-1}({M_{L}^{2n}}-{\mathbb{D}}^{2n})\oplus H_{2n-1}({M_{R}^{2n}}-{\mathbb{D}}^{2n})$
$\pi_{2n-1}^{\ast}$}
$H_{2n-1}(W^{2n}).$
@@
$[\mathbf{h}_{6n+1}\rightarrow\mathbf{h}^{\prime}_{6n+1}]=1,$
(3.2.19) $[\mathbf{h}_{6n+2}\rightarrow\mathbf{h}^{\prime}_{6n+2}]=1.$
@@
If we consider equations (3.2.7), (3.2.9), (3.2.11), (3.2), (3.2.16), (3.2.18), (3.2), (3.2), (3.2.22), and (3.2.24), then we show that the corrective term satisfies the following equation
@@ 3.3. The Reidemeister-Franz torsion of ${M^{2n}}-{\mathbb{D}^{2n}}$
$\mathbf{h}^{\prime}_{1}=\left\{\mathbf{h}^{\mathrm{Im}\,\imath_{0}^{\ast}},s_{1}(\mathbf{h}^{\mathrm{Im}\,\pi_{0}^{\ast}})\right\}$
@@
in the previous step. It follows from equation (3.3.6) that $s_{2}(\mathbf{h}^{\mathrm{Im}\,\imath_{0}^{\ast}})$ is the obtained basis $\mathbf{h}^{\prime}_{2}$ of $C_{2}(\mathcal{H}_{\ast}).$ If we take the initial basis $\mathbf{h}_{2}$ (namely, $\mathbf{h}_{0}^{\mathbb{S}^{2n-1}}$) of $C_{2}(\mathcal{H}_{\ast})$ as $s_{2}(\mathbf{h}^{\mathrm{Im}\,\imath_{0}^{\ast}})$, then we obtain
@@
By the exactness of the sequence $\mathcal{H}_{\ast}$, Lemma 3.2.1, and the First Isomorphism Theorem, we obtain the followings for each $i\in\{1,\ldots,2n-1\}$
@@
For each $i\in\{1,\ldots,2n-1\},$ by using the isomorphism
$H_{i}({M^{2n}}-{\mathbb{D}^{2n}})\overset{\pi^{\ast}_{i}}{\cong}H_{i}(M^{2n})$
and given basis $\mathbf{h}^{{M^{2n}}-{\mathbb{D}^{2n}}}_{i}$ of $H_{i}({M^{2n}}-{\mathbb{D}^{2n}}),$ we can consider the basis of $\mathbf{h}^{M^{2n}}_{i}$ of $H_{i}(M^{2n})$ as $\pi^{\ast}_{i}(\mathbf{h}^{{M^{2n}}-{\mathbb{D}^{2n}}}_{i}).$ As $\imath_{i}^{\ast}$ is zero map, $\mathrm{Im}\,\imath_{i}^{\ast}=\{0\}$. Since $\pi^{\ast}_{i}$} is an isomorphism, its inverse can be considered as the section $s_{{}_{i}}$. As in the proof of Theorem 3.2.2, we obtain
@@
$\mathbb{T}_{RF}(W_{i}^{2n},\{\mathbf{h}_{\nu}^{W_{i}^{2n}}\}_{0}^{2n})$ $=$ $\mathbb{T}_{RF}({W_{i-1}^{2n}}-{\mathbb{D}_{i}^{2n}},\{\mathbf{h}_{\nu}^{{W_{i-1}^{2n}}-{\mathbb{D}_{i}^{2n}}}\}_{0}^{2n})$
@@
and their associated Mayer-Vietoris long exact sequences as in Proposition 3.3.1. In Proposition 3.3.1, $\mathbf{h}_{\nu}^{W_{i-1}^{2n}}$ and $\mathbf{h}_{\nu}^{{M_{i+1}^{2n}}-{\mathbb{D}_{i}^{2n}}}$ are any given homology bases. So we can take these bases as above which is satisfying the equation (3.4). Since it is arbitrarily given basis in Proposition 3.3.1, we can respectively choose the same bases $\mathbf{h}_{0}^{\mathbb{S}^{2n-1}}$ and $\mathbf{h}_{2n-1}^{\mathbb{S}^{2n-1}}=\delta_{2n}(\mathbf{h}_{2n}^{W^{2n}})$ of $H_{0}(\mathbb{S}^{2n-1})$ and $H_{2n-1}(\mathbb{S}^{2n-1})$ for both sequences (3.4.3) and (3.4.4). Hence, for the basis $\mathbf{h}_{0}^{\overline{\mathbb{D}_{i}^{2n}}}=f^{i}_{\ast}(\varphi_{0}(\mathbf{c}_{0}))$ of $H_{0}(\overline{\mathbb{D}_{i}^{2n}}),$ there exist respectively bases $\mathbf{h}_{\nu}^{W_{i-1}^{2n}}$ and $\mathbf{h}_{\nu}^{M_{i+1}^{2n}}$ of $H_{\nu}({W_{i-1}^{2n}})$ and $H_{\nu}({M_{i+1}^{2n}})$ such that the following formulas hold
@@
By combining equations (3.4), (3.4.5), and (3.4.6), we obtain the Reidemeister-Franz torsion of $W_{i}^{2n}$ with untwisted $\mathbb{R}$-coefficients in these homology bases as follows
@@
4. Declarations
The author has no competing interests to declare that are relevant to the content of this article.
4.1. Ethical Approval
Research presented in this paper did not involve any sensitive data, therefore ethical approval is not necessary.
4.2. Funding
The author was partially supported by TÜBİTAK under the project number 124F247.
4.3. Availability of data and materials
Research presented in this paper has no associate data, so this declaration is not applicable.